# CPKD: Clinical Prior Knowledge-Constrained Diffusion Models for Surgical Phase Recognition in Endoscopic Submucosal Dissection

### v1

@@ 4.5 The CPKC Unit
Figure 5: (a) The order information of the phases defined in the ESD820 dataset. (b) The evaluation of an example constraint on two phases over L time steps. The constraint is first written as formula $\phi$ = ($(\lnot P7WP5)$, then represented as a DAG. (c) The computation process of $f_{0}(\lnot P7WP5,P_{t})$ is illustrated through steps (1) - (5). (1) shows the phase predictions $P_{t}$ are generated by the task model, with truth values represented visually (dark grey: True; grey: False). (2) and (3) show the evaluation of leaf nodes P7 and P5, which is the start of evaluation. (4) processes the evaluation output for P7 from (2) based on Eq.(12). In (5), the temporal logic operator W combines $\lnot P7$ and P5 following Eq.(17) to produce the complete formula evaluation
@@
Next, we define the evaluation for operators “$\neg$” (negation), $\vee$” (logical or) , and $\wedge$” (logical and):
@@ 5.1 Experiments design
Evalutation Metrics. To evaluate the performance of our surgical phase recognition model, we adopted four standard evaluation metrics widely used in prior studies [23, 24, 62, 61]: Accuracy, Precision, Recall, and Jaccard Index. Here, Accuracy measures the overall percentage of correctly classified frames in the video. To account for the class imbalance between different surgical phases, Precision, Recall, and Jaccard Index were computed per phase and then macro-averaged to obtain the final performance scores, ensuring balanced assessment across all phases.
@@ 5.2.1 Result on the ESD820 Dataset
Swin-B [56] 80.60$\pm$11.36 79.90$\pm$15.79 80.06$\pm$16.15 67.27$\pm$15.42 233
SV-RCNet [24] 82.46$\pm$14.16 84.97$\pm$17.13 83.62$\pm$18.11 72.84$\pm$20.09 30
SAHC [67] 85.01$\pm$11.92 83.22$\pm$17.00 86.44$\pm$17.27 74.03$\pm$18.93 217
Furube et al. [65] 85.63$\pm$11.36 84.43$\pm$16.24 86.48$\pm$17.76 74.94$\pm$19.01 226
AI-Endo [66] 85.24$\pm$11.10 84.45$\pm$15.02 86.05$\pm$17.04 74.72$\pm$17.92 220
SPRMamba [12] 86.35$\pm$12.59 86.04$\pm$16.35 88.80$\pm$16.62 77.90$\pm$18.11 178
ASFormer [57] 85.54$\pm$12.86 85.36$\pm$16.61 87.39$\pm$16.79 76.33$\pm$18.59 187
STFE (Ours) 87.28$\pm$9.60 87.76$\pm$12.87 88.53$\pm$12.49 79.05$\pm$15.16 224
CPKD (Ours) 88.88$\pm$10.95 89.59$\pm$14.73 91.02$\pm$14.63 82.59$\pm$16.36 218
@@ 5.2.2 Result on the Cholec80 Dataset
Our comprehensive evaluation on the Cholec80 benchmark compares the proposed approach against nine established methods: Endonet [23], MTRCNet-CL [62], LAST [30], SV-RCNet [24], TeCNO [7], TMRNet [61], Trans-SVNet [10], SAHC [67], SPRMamba [12], SR-Mamba[13], and LoViT [35]. For experimental consistency, we re-implemented TeCNO and Trans-SVNet using the authors’ provided model weights, while directly reporting published results for other comparative methods from their original papers.
Table 3: Comparison with the SOTA methods on the Cholec80 dataset. Note that the ‘+’ denotes methods based on multi-task learning that require extra tool labels and the bold text indicates the current optimal result.
Method Accuracy(%) Precision(%) Recall(%) Jaccard(%)
Endonet [23]+ 81.70$\pm$4.20 73.70$\pm$16.10 79.60$\pm$7.90 -
MTRCNet-CL [62]+ 89.20$\pm$7.60 86.90$\pm$4.30 88.00$\pm$6.90 -
LAST [30]+ 93.12$\pm$4.71 89.25$\pm$5.49 90.10$\pm$5.45 81.11$\pm$7.62
SV-RCNet [24] 85.30$\pm$7.30 80.70$\pm$7.00 83.50$\pm$7.50 -
TeCNO [7] 89.35$\pm$6.70 83.24$\pm$7.21 81.29$\pm$6.61 70.08$\pm$9.08
TMRNet [61] 90.10$\pm$7.60 90.30$\pm$3.30 89.50$\pm$5.00 79.10$\pm$5.70
Trans-SVNet [10] 90.27$\pm$6.48 85.23$\pm$6.97 82.92$\pm$6.77 72.42$\pm$8.92
SAHC [67] 91.80$\pm$8.10 90.30$\pm$6.40 90.00$\pm$6.40 81.20$\pm$5.50
SPRMamba [12] 93.12$\pm$4.58 89.26$\pm$6.69 90.12$\pm$5.61 81.43$\pm$6.90
SR-Mamba [13] 92.60$\pm$8.60 90.30$\pm$5.20 90.60$\pm$7.20 81.50$\pm$8.60
LoViT [35] 92.40$\pm$6.30 89.90$\pm$6.10 90.60$\pm$4.40 81.20$\pm$9.10
STFE+Diffusion (Ours) 93.47$\pm$6.29 89.06$\pm$4.73 91.29$\pm$7.01 82.24$\pm$9.24
The proposed method demonstrates superior performance, achieving SOTA results on multiple key metrics (Table 3). Specifically, our approach attains the highest accuracy (93.47%), surpassing the previous best-performing method, SPRMamba (93.12%). In recall, our method also leads with 91.29%, outperforming SR-Mamba (90.60%) and LoViT (90.60%). While TMRNet, SAHC, and SR-Mamba achieves a marginally higher recall (90.30%), our method remains competitive with 89.06%. Additionally, we achieve the best Jaccard index (82.24%), exceeding LoViT (81.50%) and SPRMamba (81.43%). Notably, our results are achieved without multi-task learning, unlike methods such as Endonet and LAST, which rely on additional tool annotations. This underscores the efficiency of our approach in delivering robust performance with a streamlined architecture.
@@ 5.3.2 Effectiveness of Conditioning Features
Spatial Feature 86.39$\pm$15.27 88.19$\pm$17.77 89.04$\pm$17.75 79.80$\pm$18.72
$R_{\varphi}$ Layer 7 87.20$\pm$14.90 88.89$\pm$16.93 90.55$\pm$16.74 81.65$\pm$18.23
$R_{\varphi}$ Layer 9 88.25$\pm$12.51 89.12$\pm$15.51 90.90$\pm$15.82 82.16$\pm$17.38
$R_{\varphi}$ Layer 11 88.88$\pm$10.95 89.59$\pm$14.73 91.02$\pm$14.63 82.59$\pm$16.36
$R_{\varphi}$ Layer 7,9,11 87.99$\pm$11.44 89.02$\pm$15.49 90.57$\pm$15.13 81.83$\pm$17.10
$R_{\varphi}$ Prediction 87.59$\pm$13.49 88.95$\pm$16.50 90.48$\pm$16.97 81.64$\pm$18.17

### v2

@@ 4.5 The CPKC Unit
Figure 5: (a) The order information of the phases defined in the ESD820 dataset. (b) The evaluation of an example constraint on two phases over L time steps. The constraint is first written as formula $\phi$ = $(\lnot P7WP5)$, then represented as a DAG. (c) The computation process of $f_{0}(\lnot P7WP5,P_{t})$ is illustrated through steps (1) - (5). (1) shows the phase predictions $P_{t}$ are generated by the task model, with truth values represented visually (dark grey: True; grey: False). (2) and (3) show the evaluation of leaf nodes P7 and P5, which is the start of evaluation. (4) processes the evaluation output for P7 from (2) based on Eq.(12). In (5), the temporal logic operator W combines $\lnot P7$ and P5 following Eq.(17) to produce the complete formula evaluation
@@
Next, we define the evaluation for operators $\neg$” (negation), $\vee$” (logical or) , and $\wedge$” (logical and):
@@ 5.1 Experiments design
Evalutation Metrics. Our performance evaluation employs four well-established metrics - Accuracy, Precision, Recall, and Jaccard Index. Accuracy quantifies frame-wise classification correctness across the entire video, while Precision, Recall, and Jaccard Index address class imbalance through phase-level computation followed by macro-averaging, ensuring equitable evaluation of all surgical phases regardless of their temporal prevalence. For Cholec80 benchmarking, we strictly follow the relaxed evaluation protocol [62] adopted by recent state-of-the-art methods [23, 24, 63, 61], where predictions within 10 seconds of ground truth annotations are considered valid matches. This standardized approach maintains comparability with existing research while providing a comprehensive assessment of both frame-level accuracy and phase-wise recognition quality. All other experiments utilize conventional exact-match metrics unless otherwise specified.
@@ 5.2.1 Result on the ESD820 Dataset
[…]
@@ 5.2.2 Result on the Cholec80 Dataset
Our comprehensive evaluation on the Cholec80 benchmark compares the proposed approach against nine established methods: TeCNO [7], Trans-SVNet [10], SKiT [27], LoViT [35], SPRMamba [12], SR-Mamba[13], Surgformer [69], and BNpitfalls [70]. For experimental consistency, we re-implemented TeCNO and Trans-SVNet using the authors’ provided model weights, while directly reporting published results for other comparative methods from their original papers.
Table 3: Comparison with the SOTA methods on the Cholec80 dataset.
Method Relaxed metric Accuracy(%) Precision(%) Recall(%) Jaccard(%)
[…]
As shown in Table 3, our method establishes new state-of-the-art performance across multiple metrics, achieving the highest accuracy (94.15%±6.24 vs 93.60%±5.27 for SPRMamba), recall (94.12%±3.29 vs 92.43%±6.09 for SPRMamba), and Jaccard index (85.61%±5.21 vs 84.10%±8.00 for Surgformer). While Surgformer shows a marginal advantage in precision (91.90%±4.70 vs our 91.12%±4.73), our approach maintains strong competitiveness in this metric while delivering more balanced and consistent performance across all evaluation criteria, as evidenced by lower standard deviations in most metrics. Particularly noteworthy is our method’s significant improvement in recall (1.69% absolute gain over SPRMamba) and Jaccard index (1.51% over Surgformer), which demonstrates enhanced capability in capturing complete phase segments and maintaining temporal consistency. These quantitative results validate the effectiveness of integrating diffusion models with surgical phase recognition, especially in handling complex temporal dependencies and boundary ambiguities inherent in endoscopic videos.
@@ 5.3.2 Effectiveness of Conditioning Features
[…]